ce Fand the currents I, I and I, in the curcuit shown below 1091 2.0 Ω
Answer #1

I1 + I2 = I3

loop abcda: 10 – 6I1 – 2I3 = 0

loop befcb: -4I2 – 14 + 6I1 – 10 = 0

-4I2 + 6I1 = 24

-2I2 + 3I1 = 12 ..........(1)

5 – 3I1 – I3 = 0

=> 5 – 3I1 = I1 + I2

=> 5 = 4I1 + I2

=> I2 = 5 – 4I1 ............(2)

Using the above two equations we get:

-2(5 – 4I1) + 3I1 = 12

=> -10 + 8I1 + 3I1 = 12

=> 11I1 = 22

=> I1 = 2 Amps

I2 = -3 Amps

I3 = -1 Amps

Please rate my answer if you find it helpful, good luck...